OpenStudy (anonymous):

Ravi's gas tank is 2/5 full. After he buys 8 gallons of gas, it is 2/3 full. How many gallons can Ravi's tank hold?

OpenStudy (anonymous):

what is \(\frac{2}{3}-\frac{2}{5}\) ?

OpenStudy (anonymous):

fractions are annoying, but it is possible to subtract

OpenStudy (anonymous):

ill try to elaborate more: let's call the amount of gallons that Ravi's tank can hold x and let's call the amount it has in it at the beginning y so we have: "Ravi's gas tank is 2/5 full" (2/5) * x = y now we are told that he buys 8 gallons so that means he has y+8 in it and it is 2/3 full so : (2/3) * x = y+8 we have two equations if we subtract the first from the second we get: x * ((2/3) - (2/5)) = 8 can you find x now ?